# JWST provides a new view of cosmic dawn: latest developments in studies of early galaxies

### v1

@@ Abstract
Observing and understanding the most distant light has been challenging human imagination and technical capabilities for centuries. The unique infrared capabilities of the James Webb Space Telescope (JWST) are revolutionizing the study of the most distant galaxies [e.g. Adamo24]. These data have so-far pushed the horizon at which we have observed stellar light from galaxies at redshifts [Carniani24, Naidu25], corresponding to a time of about 300 million year after the Big Bang, or a lookback time of about 13.3 billion years. A galaxy’s redshift primarily originates from the expansion of the Universe that stretches the light of a photon since its emission. The redshift corresponds to the relative scales of the Universe between emission and now as . By combining the observed redshift of the light from a galaxy with cosmological models one can derive the lookback time. JWST data have also led to the identification of new abundant populations of active galactic nuclei [Harikane23, Matthee24, Maiolino23b], they enabled unprecedented characterization of the properties of the gas, such as the chemical composition [Vanzella23, Chemerynska24, Ji24], and yielded new insights into how the light from early galaxies impacted intergalactic gas in their large-scale environments [Tang24, Witstok24, Kakiichi25, Kashino25]. Figure 1 shows zoomed-in images of JWST images of such a newly identified active galactic nucleus, one of the most primordial galaxies and the most distant galaxy known, respectively.
@@ 2 The spectrum of galaxies
The physical information that we retrieve from galaxies predominantly relies on our interpretation of their spectrum111Important exceptions to this are measurements of the clustering of galaxies, the gravitational deflection of light from background sources, and measurements of absorption lines in the light from background sources. However, all those techniques require us to detect the light from background galaxies in the first place.. The spectrum of galaxies is a composite of the light-emitting and obscuring components that include populations of stars, clouds of ionized gas around hot stars, accreting (supermassive) black holes and their surroundings, neutral gas and dust in the interstellar medium. Various detailed review articles have been written on models of the spectrum of galaxies, for example [Conroy13, Iyer25]. How can we use light to tell the difference between a young and an old galaxy? What are the key spectral features that we use to find galaxies and measure their distance? How do we know that there is an active supermassive black hole in a galaxy? What does interstellar dust do to the observed light? How can we measure detailed properties of the intergalactic gas whose emitted light itself we cannot detect?
@@ 2.2 Emission from active galactic nuclei
In addition to stellar light, a significant – in some cases dominant – source of light in distant galaxies can be an active galactic nucleus (AGN). These are luminous regions in the centers of galaxies that emit light that is powered by the accretion of matter onto a supermassive black hole (SMBH). Accreting matter forms a disk that heats up and emits emission across the electromagnetic spectrum. Moreover, processes associated with highly energetic particles give rise to nonthermal emission, such as synchrotron emission in jets, which leads to bright radio and X-Ray luminosities.
@@
Describing the X-Ray to radio spectrum of AGN requires a whole review on its own [e.g. Richards06, Siebenmorgen15]. Here, we focus on the rest-frame UV and optical spectrum of AGN. The particular emphasis of this subsection is to understand how AGN activity may mimic features of young very massive stars and, vice versa, how the presence of an AGN may be identified without usual indicators (X-Ray or radio emission), as their sensitivity is currently are currently not very sensitive in identifying faint AGN activity at . The ionizing spectrum of an AGN is typically characterised by a power-law from the accretion disk. The accretion disk has a temperature gradient such that the combined spectrum of the blackbody emission from various regions in this disk yields a power-law like spectrum.
@@ 2.4 Dust attenuation and emission
Various methods have been developed to accurately account for dust attenuation. They predominantly use the fact that dust attenuation reddens the observed spectrum, meaning that relatively more light is attenuated at shorter wavelengths (see Figure 5). For example, the reddening can be inferred by comparing the measured UV slope to an assumed intrinsic UV slope of a stellar population [Meurer99]. In SED fitting codes the reddening is included as a free parameter and the uncertainties due to degeneracies between variations in intrinsic colors and dust reddening are propagated. In some codes energy balance is assumed, meaning that there is a coupling between the attenuated UV light and the infrared emission [Cunha08]. The attenuation of the nebular light can be estimated with the Balmer decrement444In principle this method could be extended to other lines, in particular combinations of Paschen and Balmer lines that trace the same energy levels [Reddy25].. Under standard assumptions about the physical conditions of the gas, the intrinsic line-ratio of Balmer lines are known. Deviations from these intrinsic line-ratios are assumed to be due to differences in the attenuation at the respective wavelength of various transitions, and therefore be used to estimate the total attenuation.
@@ 3.2 High-redshift galaxy selections before JWST
The Lyman break technique, first used successfully for large galaxy samples at by Steidel et al. in 1996 [Steidel96] but already discussed at least two decades earlier [e.g. PartridgePeebles67, Sandage73, Koo85], relies on strong discontinuities in galaxy spectra around the Lyman- line at Å and around the Lyman-Continuum limit (912 Å), see Figure 2. These discontinuities originate from the partial ionisation state of stellar atmospheres. At high-redshift, in particular , neutral gas in the intergalactic medium additionally absorbs Lyman-Continuum and Ly photons [Madau95, Inoue14]. This effectively leads to a single (even stronger) discontinuity at the Lyman- wavelength for galaxies and quasars at redshifts beyond due to the Gunn-Peterson trough (see for example the quasar spectrum in Figure 6). The Gunn-Peterson trough is an extreme (saturated) case of similar Lyman- absorption in the regime where most gas in the intergalactic medium is neutral, which was the case before the end of cosmic reionization (see also Scheuer 1965, [Scheuer65]). This causes high-redshift galaxies to ‘drop-out’ in images taken in filters that cover rest-frame wavelengths below 1216 Å (see for example Figure 8), which happens in optical wavelengths at and at higher redshifts in the near infrared. Such galaxy samples (Lyman Break Galaxies) at specific redshift ranges are usually defined by a combination of a single or multiple colour criteria, in addition to a (continuum) magnitude limit, or by evaluating SED templates at various redshifts in a likelihood analysis. Deep optical and near-infrared imaging surveys enabled the construction of large samples at [e.g. Bouwens15, Harikane22]. The increase in 1-5 micron sensitivity enabled by JWST has enabled the Lyman break technique to successfully identify galaxies out to redshifts [Carniani24, Naidu25]. The main classes of interlopers to this technique are objects that show similarly strong colours, such as strongly dust-obscured and/or passive galaxies with a Balmer break (out to in samples with JWST; [Naidu22, ArrabalHaro23], and brown dwarf stars [e.g. Bowler15, GRB24]). These interlopers are particularly an issue for luminous high-redshift galaxies that are often identified in fields with shallower data or sparser wavelength coverage, and less of an issue for fainter systems as the number counts of high-redshift galaxies are typically much steeper than those of interlopers, such that the likelihood of encountering interlopers decreases towards fainter luminosities.
@@
The most commonly used emission-line selection identifies galaxies based on their Lyman- emission (hence Lyman- emitters) and also dates back to the 90s [e.g. Hu96]. The Lyman- line is intrinsically one of the brightest emission-lines produced in HII regions, and its rest-frame UV wavelength makes it very suitable for identifying galaxies at redshifts . Indeed, the narrow-band technique, where emission-line galaxies are identified based on a colour excess over the photometry in an overlapping broad-band has yielded samples of thousands of Lyman- selected galaxies at [e.g. Sobral18, Ouchi20]. The main contaminants for tese samples are other emission-lines from lower redshift galaxies, primarily [Oiii] and H. Additional photometric information is used to identify such interlopers [e.g. Matthee17]. With the advent of wide-field integral field spectroscopy, in particular thanks to the Multi Unit Spectroscopic Explorer on the Very Large Telescope, direct spectroscopic high-redshift Lyman- selection has been very efficient [Bacon17]. Due to the high spectral slicing ( Å), the contrast of narrow emission lines over the sky background is very high, yielding unprecedented (surface) brightness limits, enabling the detection of diffuse light [Wisotzki18] and galaxies that are undetected even in the deepest HST data [Maseda18].
@@ 3.3 High-redshift galaxy selections in the JWST era
Since the summer of 2022, the JWST has made immense improvements in our ability to identify and study distant galaxies. This is enabled by various key technological advances thanks to JWST’s large passively cooled 6m mirror and the diverse instrument suite. The Near-InfraRed Camera (NIRCam) enables sensitive imaging in various filters covering the m regime at resolution and wide-field slitless spectroscopy over 3-5 m. The Mid-Infrared Instrument (MIRI) is optimised for imaging and spectroscopy over 5-30 m at resolution. The Near Infrared Spectrograph (NIRSpec) enables single and multi-object spectroscopy over the 1-5 m range, as well as an integral field mode, with a spectral resolution from 100 - 3000 km s-1. Moreover, the Near Infrared Imager and Slitless Spectrograph (NIRISS) enables highly efficient wide-field slitless spectroscopy over m and imaging from 1-5 m. The field of view of JWSTs instrument ranges from 3 arcsec2 (NIRspec and MIRI integral field spectroscopy) to arcmin2 (NIRcam imaging). Among these cameras that all significantly improve upon previous available instruments, the most significant gain has been made in the spectroscopic capabilities over the 2-5 m wavelength regime, where the sensitivity improvement has been more than three orders of magnitude, in addition to the new multiplexing capabilities.
@@ 4 Recent developments, open questions and controversies
In this final part of the review, we aim to provide an overview of research topics that have been particularly actively studied, emphasizing the main objectives and new methods and datasets that are driving progress and what are currently some of the main limitations. This list of topics is certainly incomplete and biased by my personal work and background, and barely touches upon areas such as studies of galaxy structures, kinematics, spatially resolved properties or the properties of dust (see [Adamo24, Stark25, Ellis25] for other overview of recent results on these topics).
@@ 4.1 Early star formation
The history of star formation in the Universe is among the most fundamental properties that observational astronomers aim to infer. One of the main advances of the JWST is that it enables us to obtain more and better probes of galaxies’ star formation activity beyond redshifts . Before JWST, the most sensitive constraints on the evolution of the cosmic star formation rate density primarily based on rest-frame UV data, taken, for example, with the Hubble Space Telescope [e.g. Bouwens15]. These measurements were affected by two main limitations. First, the observed UV light from young stars may be strongly attenuated by dust in the interstellar medium of galaxies up to degrees that samples of strongly obscured galaxies may have been missing from the samples. Direct measurements of the density of obscured star formation have been made based on infrared and sub-millimeter data [Gruppioni13, HodgeCunha20], but these have typically not reached sensitivities comparable to rest-frame UV data. Second, the redshifts of the vast majority of galaxy samples at were estimated based on photometry, which may contain biases or be subject to interlopers. With JWST these challenges can be addressed thanks to the large number of spectroscopic redshifts and the possibility to select galaxies by their H line emission (a sensitive tracer of ongoing star formation that is significantly less subject to dust attenuation; [Sun25, diCesare25]). Moreover, JWST data have advanced the redshift frontier (), yielded new insights into the star formation histories of individual reionization-era galaxies, and probed the assembly of the first massive galaxies.
@@ 4.1.1 A bright cosmic dawn
The number of (luminous) galaxies discovered by JWST at redshifts beyond have defied expectations [Naidu22, Castellano22, Adams23]. Although the data-sets that reach into new regions of parameter space have also suffered from new classes of interlopers (such as dust-obscured emission-line galaxies at redshifts , [Naidu22, ArrabalHaro23]), multiple studies reported a factor of a few more luminous galaxies at [Bouwens23, Donnan24, Whitler25], compared to the expectations (see Figure 10). These expectations were based on models that assume a constant efficiency of star formation from gas accreting on dark matter halos, whose evolution can be predicted for a given cosmological model, and a fixed efficiency of UV light production per unit star formation rate [Mason15, Tacchella18]. These models successfully matched the evolution of the UV luminosity density after the first Gyr of cosmic. What could be causing these unexpectedly high numbers of earliest galaxies? While there are models that accelerate structure formation itself (effectively changing our cosmological model, see [MKB23] for a discussion), the majority of possibly explanations are astrophysical. One suggestion is that the star formation efficiency was higher in the early Universe, for example, because supernova explosions may not strongly suppress star formation (so-called supernova feedback) in early galaxies. This may happen in the high gas densities in early galaxies where the supernova feedback is delayed with respect to the timescales on which stars form [e.g. Dekel23]. Alternatively, a more efficient conversion of stellar mass to UV light could yield a higher number of galaxies without the need to form more stellar mass. This could be achieved either by a negligible dust attenuation [Ferrara23] or by processes that change the relative fraction of massive stars that contribute to the galaxies’ spectra, such as younger ages [Whitler23], ‘bursty’ star formation rates that rapidly vary [GSun23], an evolving contribution of nebular continuum emission [Katz24] and/or a top heavy IMF that has relatively more massive stars [Cueto24]. Finally, it is also possible that there are additional sources of UV light that may become important at higher redshift, such as AGNs. Several of these scenarios can be addressed directly with more spectroscopic follow-up observations, but also by testing models for their consistency with redshift evolution of the shape of the UV luminosity function, or other observables such as the build-up of the galaxy stellar mass function [e.g. Donnan24].

### v2

@@ Abstract
Observing and understanding the most distant light has been challenging human imagination and technical capabilities for centuries. Now, the unique infrared capabilities of the James Webb Space Telescope (JWST) are revolutionizing the study of the most distant galaxies [e.g. Adamo24]. Data from JWST have so-far allowed us to push the horizon at which we have observed stellar light from galaxies at redshifts [Carniani24, Naidu25], corresponding to a time of about 300 million year after the Big Bang, or a lookback time of about 13.3 billion years. A galaxy’s redshift primarily originates from the expansion of the Universe that stretches the light of a photon since its emission. The redshift corresponds to the relative scales of the Universe between emission and now as . By combining the observed redshift of the light from a galaxy with cosmological models one can derive the lookback time. JWST data have also led to the identification of new abundant populations of active galactic nuclei [Harikane23, Matthee24, Maiolino23b], they enabled unprecedented characterization of the properties of interstellar gas, such as the chemical composition [Vanzella23, Chemerynska24, Ji24], and yielded new insights into how the light from early galaxies impacted the intergalactic medium in their large-scale environments [Tang24, Witstok24, Kakiichi25, Kashino25]. Figure 1 shows zoomed-in images of JWST images of such a newly identified active galactic nucleus, one of the most primordial galaxies and the most distant galaxy known, respectively.
@@ 2 The spectrum of galaxies
The physical information that we retrieve from galaxies predominantly relies on our interpretation of their spectrum111Important exceptions to this are measurements of the clustering of galaxies, the gravitational deflection of light from background sources, and measurements of absorption lines in the light from background sources. However, all those techniques require us to detect the light from background galaxies in the first place.. The spectrum of galaxies is a composite of the light-emitting and obscuring components that include populations of stars, clouds of ionized gas around hot stars, accreting (supermassive) black holes and their surroundings and neutral gas and dust in the interstellar medium. Various detailed review articles have been written on models of the spectrum of galaxies, for example [Conroy13, Iyer25]. How can we use light to tell the difference between a young and an old galaxy? What are the key spectral features that we use to find galaxies and measure their distance? How do we know that there is an active supermassive black hole in a galaxy? What does interstellar dust do to the observed light? How can we measure detailed properties of the intergalactic gas whose emitted light itself we cannot detect?
@@ 2.2 Emission from active galactic nuclei
In addition to stellar light, a significant – in some cases dominant – source of light in distant galaxies can be an active galactic nucleus (AGN). This is a luminous region in the center of a galaxy that emits light that is powered by the accretion of matter onto a supermassive black hole (SMBH). Accreting matter forms a disk that heats up and emits emission across the electromagnetic spectrum. Moreover, processes associated with highly energetic particles give rise to nonthermal emission, such as synchrotron emission in jets, which leads to bright radio and X-Ray luminosities.
@@
Describing the X-Ray to radio spectrum of AGN requires a whole review on its own [e.g. Richards06, Siebenmorgen15]. Here, we focus on the rest-frame UV and optical spectrum of AGNs. The particular emphasis of this subsection is to understand how AGN activity may mimic features of young very massive stars and, vice versa, how the presence of an AGN may be identified without usual indicators (X-Ray or radio emission), as their sensitivity is currently not very constraining in identifying faint AGN activity at . The ionizing spectrum of an AGN is typically characterised by a power-law from the accretion disk. The accretion disk has a temperature gradient such that the combined spectrum of the blackbody emission from various regions in this disk yields a power-law like spectrum.
@@ 2.4 Dust attenuation and emission
Various methods have been developed to accurately account for dust attenuation. They predominantly use the fact that dust attenuation reddens the observed spectrum, meaning that relatively more light is attenuated at shorter wavelengths (see Figure 5). For example, the reddening can be inferred by comparing the measured UV slope to an assumed intrinsic UV slope of a stellar population [Meurer99]. In SED fitting codes the reddening is included as a free parameter and the uncertainties due to degeneracies between variations in intrinsic colors and dust reddening are propagated. In some codes energy balance is assumed, meaning that there is a coupling between the attenuated UV light and the infrared emission [Cunha08]. The attenuation of the nebular light can be estimated with the Balmer decrement444In principle this method could be extended to other lines, in particular combinations of Paschen and Balmer lines that trace the same energy levels [Reddy25].. Under standard assumptions about the physical conditions of the gas, the intrinsic line-ratio of Balmer lines are known. Deviations from these intrinsic line-ratios are assumed to be due to differences in the attenuation at the respective wavelength of various transitions, and therefore can be used to estimate the total attenuation.
@@ 3.2 High-redshift galaxy selections before JWST
The Lyman break technique, first used successfully for large galaxy samples at by Steidel et al. in 1996 [Steidel96], but already discussed at least two decades earlier [e.g. PartridgePeebles67, Sandage73, Koo85], relies on strong discontinuities in galaxy spectra around the Lyman- line at Å and around the Lyman-Continuum limit (912 Å), see Figure 2. These discontinuities originate from the partial ionisation state of stellar atmospheres. At high-redshift, in particular , neutral gas in the intergalactic medium additionally absorbs Lyman-Continuum and Ly photons [Madau95, Inoue14]. This effectively leads to a single (even stronger) discontinuity at the Lyman- wavelength for galaxies and quasars at redshifts beyond due to the Gunn-Peterson trough (see for example the quasar spectrum in Figure 6). The Gunn-Peterson trough is an extreme (saturated) case of similar Lyman- absorption in the regime where most gas in the intergalactic medium is neutral, which was the case before the end of cosmic reionization (see also Scheuer 1965, [Scheuer65]). This causes high-redshift galaxies to ‘drop-out’ in images taken in filters that cover rest-frame wavelengths below 1216 Å (see for example Figure 8), which happens in optical wavelengths at and at higher redshifts in the near infrared. Such galaxy samples (Lyman Break Galaxies) at specific redshift ranges are usually defined by a combination of a single or multiple colour criteria, in addition to a (continuum) magnitude limit, or by evaluating SED templates at various redshifts in a likelihood analysis. Deep optical and near-infrared imaging surveys enabled the construction of large samples at [e.g. Bouwens15, Harikane22]. The increase in 1-5 micron sensitivity enabled by JWST has enabled the Lyman break technique to successfully identify galaxies out to redshifts [Carniani24, Naidu25]. The main classes of interlopers to this technique are objects that show similarly strong colours, such as strongly dust-obscured and/or passive galaxies with a Balmer break (out to in samples with JWST; [Naidu22, ArrabalHaro23], and brown dwarf stars [e.g. Bowler15, GRB24]). These interlopers are particularly an issue for luminous high-redshift galaxies that are often identified in fields with shallower data or sparser wavelength coverage, and less of an issue for fainter systems as the number counts of high-redshift galaxies are typically much steeper than those of interlopers, such that the likelihood of encountering interlopers decreases towards fainter luminosities.
@@
The most commonly used emission-line selection identifies galaxies based on their Lyman- emission (hence Lyman- emitters) and also dates back to the 90s [e.g. Hu96]. The Lyman- line is intrinsically one of the brightest emission-lines produced in Hii regions, and its rest-frame UV wavelength makes it very suitable for identifying galaxies at redshifts . Indeed, the narrow-band technique, where emission-line galaxies are identified based on a colour excess over the photometry in an overlapping broad-band has yielded samples of thousands of Lyman- selected galaxies at [e.g. Sobral18, Ouchi20]. The main contaminants for these samples are other emission-lines from lower redshift galaxies, primarily [Oiii] and H. Additional photometric information is used to identify such interlopers [e.g. Matthee17]. With the advent of wide-field integral field spectroscopy, in particular thanks to the Multi Unit Spectroscopic Explorer on the Very Large Telescope, direct spectroscopic high-redshift Lyman- selection has been very efficient [Bacon17]. Due to the high spectral slicing ( Å), the contrast of narrow emission lines over the sky background is very high, yielding unprecedented (surface) brightness limits, enabling the detection of diffuse light [Wisotzki18] and galaxies that are undetected even in the deepest HST data [Maseda18].
@@ 3.3 High-redshift galaxy selections in the JWST era
Since the summer of 2022, the JWST has made immense improvements in our ability to identify and study distant galaxies. This is enabled by various key technological advances thanks to JWST’s large passively cooled 6m mirror and the diverse instrument suite. The Near-InfraRed Camera (NIRCam) enables sensitive imaging in various filters covering the m regime at resolution and wide-field slitless spectroscopy over 3-5 m. The Mid-Infrared Instrument (MIRI) is optimised for imaging and spectroscopy over 5-30 m at resolution. The Near Infrared Spectrograph (NIRSpec) enables single and multi-object spectroscopy over the 1-5 m range, as well as an integral field mode, with a spectral resolution from 100 - 3000 km s-1. Moreover, the Near Infrared Imager and Slitless Spectrograph (NIRISS) enables highly efficient wide-field slitless spectroscopy over m and imaging from 1-5 m. The field of view of JWST’s instrument ranges from 9 arcsec2 (NIRspec and MIRI integral field spectroscopy) to arcmin2 (NIRCam imaging). Among these cameras that all significantly improve upon previous available instruments, the most significant gain has been made in the spectroscopic capabilities over the 2-5 m wavelength regime, where the sensitivity improvement has been more than three orders of magnitude, in addition to the new multiplexing capabilities.
@@ 4 Recent developments, open questions and controversies
In this final part of the review, we aim to provide an overview of research topics that have been particularly actively studied, emphasizing the main objectives and new methods and datasets that are driving progress and what are currently some of the main limitations. This list of topics is certainly incomplete and biased by the author’s personal work and background, and barely touches upon areas such as studies of galaxy structures, kinematics, spatially resolved properties or the properties of dust (see [Adamo24, Stark25, Ellis25] for other overviews of recent results on these topics).
@@ 4.1 Early star formation
The history of star formation in the Universe is among the most fundamental properties that observational astronomers aim to infer. One of the main advances of the JWST is that it enables us to obtain more and better probes of galaxies’ star formation activity beyond redshifts . Before JWST, the most sensitive constraints on the evolution of the cosmic star formation rate density were primarily based on rest-frame UV data, taken, for example, with the Hubble Space Telescope [e.g. Bouwens15]. These measurements were affected by two main limitations. First, the observed UV light from young stars may be strongly attenuated by dust in the interstellar medium of galaxies up to levels that samples of strongly obscured galaxies may have been missing from the samples. Direct measurements of the obscured star formation rate density have been made based on infrared and sub-millimeter data [Gruppioni13, HodgeCunha20], but these have typically not reached sensitivities comparable to rest-frame UV data. Second, the redshifts of the vast majority of galaxy samples at were estimated based on photometry, which may contain biases or be subject to interlopers. With JWST these challenges can be addressed thanks to the large number of spectroscopic redshifts and the possibility to select galaxies by their H line emission (a sensitive tracer of ongoing star formation that is significantly less subject to dust attenuation; [Sun25, diCesare25]). Moreover, JWST data have advanced the redshift frontier (), yielded new insights into the star formation histories of individual reionization-era galaxies, and probed the assembly of the first massive galaxies.
@@ 4.1.1 A bright cosmic dawn
The number of (luminous) galaxies discovered by JWST at redshifts beyond have defied expectations [Naidu22, Castellano22, Adams23]. Although the data-sets that reach into new regions of parameter space have also suffered from new classes of interlopers (such as dust-obscured emission-line galaxies at redshifts , [Naidu22, ArrabalHaro23]), multiple studies reported a factor of a few more luminous galaxies at [Bouwens23, Donnan24, Whitler25], compared to the expectations (see Figure 10). These expectations were based on models that assume a constant efficiency of star formation from gas accreting on dark matter halos, whose evolution can be predicted for a given cosmological model, and a fixed efficiency of UV light production per unit star formation rate [Mason15, Tacchella18]. These models successfully matched the evolution of the UV luminosity density after the first Gyr of cosmic time. What could be causing these unexpectedly high numbers of early galaxies? While there are models that accelerate structure formation itself (effectively changing our cosmological model, see [MKB23] for a discussion), the majority of possible explanations are astrophysical. One suggestion is that the star formation efficiency was higher in the early Universe, for example, because supernova explosions may not strongly suppress star formation (so-called supernova feedback) in early galaxies. This may happen in the high gas densities in early galaxies, where the supernova feedback is delayed with respect to the timescales on which stars form [e.g. Dekel23]. Alternatively, a more efficient conversion of stellar mass to UV light could yield a higher number of galaxies without the need to form more stellar mass. This could be achieved either by a negligible dust attenuation [Ferrara23] or by processes that change the relative fraction of massive stars that contribute to the galaxies’ spectra, such as younger ages [Whitler23], ‘bursty’ star formation rates that rapidly vary [GSun23], an evolving contribution of nebular continuum emission [Katz24] and/or a top heavy IMF that has relatively more massive stars [Cueto24]. Finally, it is also possible that there are additional sources of UV light that may become important at higher redshift, such as AGNs. Several of these scenarios can be addressed directly with more spectroscopic follow-up observations, but also by testing models for their consistency with redshift evolution of the shape of the UV luminosity function, or other observables such as the build-up of the galaxy stellar mass function [e.g. Donnan24].